INAF queue observing, 2016 Feb 29- Mar 1

Observer: R. Carini, M. Faccini
Telescope Operator: G. Bechetti
Support Astronomer: O. Kuhn
Instrument: LBC, LUCI

Summary

At the sunset the sky is clear

Night duration (12\xB0 el twilight) 19:30-5:30 ( 02:30-12:30 UT) = 10 h 00m

Observing time: 8h 30m (85%)

Weather time loss: 0 h ()

technical time loss 1 h 30 m (15%)

Data Summary

Project_RUN Objects Filter ExpTime Notes
Flat LBC blue: V, B; red: V, r    
PHOT std S95 LBC-b(U_spec, V,B), LBC-r( V,r,i,I)    
Marconi IZw18 LBC-b(V); LBC-r(I) 10 m 7th ob done, seeing out of specific
Dotto 2007 VA85, 313276 blue: V,B,g; red: R,z,r,I 8 m completed
Raimondo UmaII LBC-b(U_spec, B); LBC-r( r) 1h 12 m field 6 is good, to be repeated the red of the 7th and all 8th field (seeing too high)
Dotto 426071, 267221 blue: V,B,g; red: R,z,r,I 8 m completed
Bellazzini DDT Secco_C1 dual grating, r sloan 40 m to be repeated
standard Hz44 dual grating    

Details

01:29 start the flat field

02:08 photometric standard SA95
  • dofpia
  • 02:17 run the ob
  • seeing 1.2'''

02:27 slew to Marconi's program
  • run IZw18_copointing.ob
  • dofpia
  • lbcrangebal
  • 02:41 run science IZw18_6.ob
  • seeing 1.4''-1.5''

03:08 the seeing is not so good, we slew to Dotto's program
  • target V2007 VA85
  • run copointing
  • dofpia
  • lbcrangebal
  • 03:15 run the scientific ob 2007VA85a.ob and set the ephimeres
  • seeing 1.1''-1.2'''
  • 03:25 run 2007VA85b.ob

3:28 slew to the target 313276
  • copointing
  • preset failed, problem with the copointing
  • new copointing
  • dofpia
  • lbcrangebal
  • 3:51 run 313276.ob
  • seeing 1''-1.1''

4:00 the seeing seems stable around 1''-1.1'', slew to Raimondo program field 6
  • copointing UMaII _f6_UBr_point.ob
  • dofpia
  • lbcrangebal
  • 4:14 run UMaII _f6_UBr.ob
  • 4:22 mean seeing 0.8''-0.9'''
  • 4.30 mean seeing 1.1''-1.2'''

5:04 run dofpia for UMaII _f7_UBr.ob
  • 5:08 run UMaII _f7_UBr.ob
  • seeing 0.8''-1.0''
  • 5:33 seeing 0.85''
  • 5:40 seeing 0.85''-0.95''
  • 5:51 seeing 0.9''
  • 5:58 seeing 1.2''

06:02 dofpia for UMaII _f8_UBr.ob
  • 06:06 run UMaII _f8_UBr.ob
  • 06:40 the seeing is too much variable and high 1.2''-1.5''. stop the acquisition of the data. this field have to be repeated

06:40 slew to Dotto's program, target: 426071
  • 06:43 run copointing, dofpia, lbcrangebal
  • run the ob
  • 7:02 problem with the readout in the red channel( chip 4). the images from lbcr.20160309.050909 in red channel are not good. To better understand the problem, see the technical issue
send by the SA.

07:03 disconnect and connect LBC

07:10 run again the copointg fot the target 426071
  • dofpia, lbcrangebal
  • 07:22 run the scientific target 426071.ob

07:30 slew to the target 267221
  • dofpia and lbcrangebal
  • 07:45 problems with Rb_science and iraf
  • try again the lbcrangebal, but there are again problems
  • 08:00 lbcrangebal done and run the scientific ob 267221.ob

8:05 switch to LUCI2

08:25 in the field of the telluric star there is not a star to collimate, the TO have to use a star in another part of the sky

08:50 collimation done * seeing about 0.9''-1.0''
  • load script of the telluric star for the program of deBarros

08:55 error with the mask, while trying to move the mask from the focal plane unit to the turnout position at MOS
  • Dave Thompson are trying to fix the problem

09:50 the problem can not be fixed in a few time, we switch to Bellazzini program in parallel mode (LBC-r + MODS-dual)
  • the TO configures the instruments

09:59 wake up MODS
  • sieveSnap.pro mods1r.20160301.0001,mods1b.20160301.0001

10:30 acqMODS SECCO_C1.acq mods1r.0002-0004
  • 10:33 dofpia
  • the guider can not guide
  • 10:58 we change the guide star. acqMODS SECCO_C1.acq mods1r.0005-0008
  • dofpia

11:13 execMODS SECCO_C1.obs =mods1r.0009-00010,mods1b.0002-0003 =
  • seeing 1.1''-1.3''
  • run ob
  • 12:03 we stop the science (the data are not in the constraints)

12:05 acquisition standard star acqMODS Hz44.acq mods1r.0011-00014
  • 12:42 execMODS Hz44.obs =mods1r.0015-00016,mods1b.0004-0005 =

12:42 Flat field with LBC -- %USERSIG{RobertaCarini - 2016-03-01}%

Comments

 
Topic revision: r9 - 03 Mar 2016, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback